OpenStudy (anonymous):

What are the steps for multiplying monomial algebraic terms where they have exponents with like bases?

OpenStudy (anonymous):

can you show me an example?

OpenStudy (anonymous):

The area, 4x2, is a product of a number (4) and a variable with a whole number exponent (x2). In other words, it's a monomial, too. So the result of multiplying two monomials is—another monomial! Let's try a slightly more complicated problem. Let's find the area of a circle with a radius of 2xy. The formula for the area of a circle is A = pr2, where A= area and r = radius. To find the area of a circle with a radius of 2xy, we'll first need to square the radius, and then multiply by p.

OpenStudy (anonymous):

Like in a^x * a^y = a*(x+y) ?

OpenStudy (anonymous):

I'm not sure..

OpenStudy (anonymous):

you are right. you can be sure

OpenStudy (anonymous):

I noticed the attachment to your last post. In that example you calculate the area of a circle with radius 2xy via: PI * (2xy)^2 = PI * 2 * X * Y * @ * X * Y, simplified into: PI * 4 * x^2 * y^2 Is that the result you are looking for.

OpenStudy (anonymous):

Thanks guys !

OpenStudy (anonymous):

& Yes !